OpenStudy (anonymous):

Which of the following is one of the factors of 2x2 + 2x – 144? Answer x + 9 x – 8 2 All of the above

OpenStudy (anonymous):

all of the above 2x^2 + 2x – 144 = 2(x^2 + x - 72) = 2(x + 9)(x - 8)